General Meeting Summary 4/17/14
NPS PROTOTYPE STATUS:
- HV dividers: Mistake with divider found and corrected. Divider assemblies are being redone by the electronics group. Also received R4125 and divider from FIU. Should have enough PMTs and dividers for prototype tests now.
- LED testing: Intensity measurements with LED at different distances from the photodiode (S2281)
- Output of IR LEDs is linear with driving current up to about 60 mA and then saturates. The saturation does not seem to be due to the photodiode. This was checked with measurements with and without filter, which gave the same result. Response of blue LEDs is similar to that of the IR LEDs.
- PMT sensitivity to IR light was tested with the LED (peak >950nm) at different distances from the PMT.
- Results suggest that PMT is still sensitive. Additional tests were done with a PbWO4 or PbF2 crystal inserted between the PMT and the LED and it was found that the PMT is less sensitive in this configuration. One thought about the PMT sensitivity was that the LED emits not only in the IR but also visible region. Tests were repeated with a long pass filter cutting all wavelengths but IR, but the PMT sensitivity remained. This suggests that the majority of the light has IR wavelengths and the PMT could receive long term damage if left on during IR curing.
- Further testing will include measurements with LEDs of peak wavelength >1000 nm and better filters to protect the PMT. Discussion about trying to quantify how much one shortens the PMT lifetime with IR curing by, e.g., analyzing the relation betwen LED driving current and PMT anode current. In the future one could also try to do in-situ studies with beam.
- Crystals: Transmittance measurements of PbWO4 and PbF2 crystals
- Results show that the transmittance for both crystals saturates above 400 nm. So using a common light source with wave length ~>500 nm for a Light Monitoring System with a combined PbWO4/PbF2 crystal matrix should be fine.
- Radiation studies in which transmittance will be tested before and after irradiation of the crystal will begin as soon as blue/IR curing systems are constructed
DISCUSSION ABOUT ELECTRONICS TESTS
- Setup mostly ready and will start testing soon. Some additional hardware was recently purchased to alleviate scheduling issues.
CALOR 2014 - PWO DISCUSSION
- R. Novotny reported that his recent experience with SICCAS had not been very good. Based on tests of a recent batch of ~370 crystals it was found that 30-40% had to be rejected. Replacements were delivered, but the quality was not improved.
- Based on the discussion it seems that R. Novotny now prefers to work with Crytur on the remaining ~8000 crystals needed for the EM calorimeter. He has already set up a contract to work closely with the company on the development and production of the crystals.
- Discussion about Crytur vs. Siccas. Overall it seems that it may be beneficial to work with a company like Crytur that is interested in doing R&D. Crytur is already known in the PWO field. They are using the same crystal growth method as BTCP though may not have as much experience with the doping process, which is important for light yield and radiation hardness. Apparently, Crytur had been working with CMS on PWO crystal development, but were then rejected over BTCP for the final CMS procurements. The fact that Crytur is based in Europe may have additional benefits.
- There is an opportunity for NPS to closely collaborate with Panda on the effort with Crytur. It would be good to check the PWO requirements of NPS and Panda again.
- Discussion about submitting a proposal to the EIC R&D opportunity. The idea would be to set up a testing facility for Crytur produced crystals, which could be used for the planned EIC calorimeter, which is envisioned to be PWO based, and test different methods of readout. Next steps will include contacting C. Woody and set up a meeting with G. Young to discuss about the PWO supply problems and possible solution.
ITEMS FOR NEXT MEETING:
- NPS design and prototype
- Electronics tests
- RCS/pi0 proposal status
NEXT MEETING: THUR, 8 MAY at 9AM (EST)